# Inference of time delay in stochastic systems

### v1

@@ IV.4 General nonlinear time-delayed feedback
However, we can make use of the observation for the linear system that the minima and maxima are $\Delta\omega\approx 2\pi/\tau$ apart in frequency space for large enough values of $\omega$. We hypothesize that oscillations with this oscillation period can be found in the $\omega^{2}$ of nonlinear systems as well, i.e., that the signature of time delay carries over from linear to more general nonlinear systems. On this basis, we can design a numerical procedure based on a polynomial fit in order to extract the positions of the minima and maxima in frequency space from numerical data and subsequently compute an approximate inferred delay time (for details of the procedure, see Appendix A.3).
@@
Figure 6: Numerically simulated PSD (red) as a function of $\omega$ for non-linear systems (12), (Panel a) (13) (Panel b), and  (14) (Panel c). In each panel: Polynomial fit (blue) using the quartic splines and automatically extracted positions (cyan circles) of maxima and minima using the roots of the derivative of the polynomial (see Appendix A.3). See Table 1 for the corresponding inference results.
@@
Figure 6 presents exemplary fits of the resulting function $\omega^{2}$ including the obtained positions of the minima and maxima, that were used to infer the delay time. The extracted locations of the minima and maxima are in good agreement with the data as shown in Fig. 6, where the extracted locations are marked as circles.
@@
$0.1$ $0.101\pm 0.008$ $0.1$ $0.101\pm 0.003$ $0.1$ $0.097\pm 0.007$
$0.15$ $0.151\pm 0.003$ $0.15$ $0.152\pm 0.003$ $0.15$ $0.150\pm 0.002$
$0.3$ $0.303\pm 0.012$ $0.3$ $0.303\pm 0.007$ $0.3$ $0.319\pm 0.027$
Table 1: Delay times $\tau_{\mathrm{inferred}}$ inferred from PSD with one standard deviation as uncertainty: (a) Data corresponding to Eq. (12) with $a=10$ and $b=c=10$; (b) Data corresponding to Eq. (13) with $a=b=10$ and $c=20$; Data corresponding to Eq. (14) with $a=10$, $b=5$ and $c=20$.
Overall, we find robust and accurate inference results for various delay times with an uncertainty of the order of one percent or less. The accuracy of the inference is similar for the different linear and nonlinear feedback and instantaneous forces. Crucially, for the described procedure, knowing the functional form of the forces (instantaneous and delayed) is not required. Therefore, this procedure can be applied to a wide range of systems with discrete time delay.
@@ V Perturbation and deep learning-based approach – Methodology
Up to this point, we have demonstrated that the power spectral density (PSD) features a characteristic signature of time delay for a wide range of feedback forces (linear and nonlinear) in stochastic systems. This allowed us to infer time delay even if the exact functional form of the underlying delayed and instantaneous forces is unknown. As a drawback, this methodology requires a large number of long trajectories. This can be a challenge in experiments, where one may only have fewer and shorter trajectories at ones disposal.
@@
To circumvent this issue, we here provide an alternative procedure to infer time delay. This procedure is based on the observation of the response of the system under the influence of a perturbative force. Broadly speaking, for a delay-driven system subjected to a perturbation, we expect the system’s response not instantaneous but rather with some delay, as we shall see explicitly in the following Subsec. V.1. Assuming the perturbation to be a force that is applied instantaneously and that is sufficiently strong (i.e., strong compared to noise, feedback forces and instantaneous forces), we expect a distinct delayed reaction, a characteristic response, which we explain in detail in the following.
@@ V.1 Characteristic response of a system with time delay to a perturbation
To develop an analytical understanding of the characteristics of the response, we take one step back and consider a deterministic system (i.e., the limit $T\to 0$). Here, we consider a system with linear time-delayed feedback force with an additional perturbative force $F_{\mathrm{perturb}}(t)$. Its equation of motion reads:
@@
Intuitively, in the limit $k_{a}=k_{b}>0$ (attractive harmonic), the linear feedback force can be understood as the consequence of a delayed harmonic potential centered at $x(t-\tau)$. Figure 7 shows its schematic representation. We assume that the particle was at rest at $x(t)=0$ in the time interval $t\in[-\tau,0]$. We then apply a perturbative force, $F_{\rm perturb}$, starting at $t=0$ for a duration $T_{\mathrm{perturb}}>\tau$. The particle will first be confined to a harmonic potential centered at $x(t-\tau)=0$. Following dynamics (15) after $t=\tau$, the delayed position $x(t-\tau)$ (and hence the location of the harmonic trap) will also start to change, resulting in the abrupt change in dynamical behavior.
@@
Figure 8: Effect of a perturbation on deterministic (a-c) and stochastic systems (d-e). Panels (a-c): The particle subject to attractive harmonic time-delayed feedback and external perturbation (15). Red dashed lines: Analytical solutions for position (17) [panel (a)] and velocity (18) [panel (b)]. Dash-dotted vertical line indicates end of perturbation; dashed grey vertical lines represent the beginning and end of observation; vertical dotted lines (color-coded) serve as a guide for the eye at one delay time $\tau$ after the onset, and then, after the end of the perturbation. Panel (d) shows discretized ‘velocity’ $v_{x,\mathrm{norm}}$ after noise averaging over $25$ realizations, applying a simple moving average over three time steps and normalization with the maximum velocity $v_{\mathrm{max}}$ as a function of time during the observation window; the deterministic velocity (black, from panel b) is shown as a reference. Dotted vertical lines serve as a guide for the eye at $t=\tau$ and $t=T_{\mathrm{perturb}}+\tau$. Panels (c and e): The perturbing force $F_{\mathrm{perturb}}$ as a function of time. The system is perturbed for a duration of $T_{\mathrm{perturb}}=0.25>\tau$ (a-c) and $T_{\mathrm{perturb}}=0.25\tau_{B}>\tau$ (d-e), and the observation continues for another $T_{\mathrm{perturb}}$.
@@
To make analytical progress, we now assume a perturbation protocol given by a step function, i.e., $F_{\mathrm{perturb}}(t)=F\Theta(t)$, where $\Theta(t)$ is the Heaviside function. We calculate the dynamics resulting from Eq. (15) using the so-called method of steps [66]. This method can be used to obtain an analytical solution in intervals of one delay time iteratively, provided that the history function $\Phi(t)$ containing the position of the particles in the previous time interval is known, i.e., $\Phi(t)=x(t)$ for $t\in[t_{0}-\tau,t_{0}]$. Formally, its piecewise solution as a function of time during the interval $[t_{0},t_{0}+\tau]$ reads
[…]
@@
Similarly, the particle’s position $x(t)$ can be evaluated for $t\geq t_{0}+\tau$ iteratively. However, for our analytical considerations, we focus on the dynamical behavior during first two $\tau$ after switching on the perturbation at $t=0$.
@@
In our case, the initial history function represents a particle that is at rest in position $x=0$, i.e., we choose $\Phi(t)=x(t)=0$ for $t\in[-\tau,0]$ (which can easily be realized by pinning the particle). The analytical expressions for the position as a function of time during the time interval $[-\tau,2\tau]$ following Eq. (16) read as follows:
@@
For reasons that will become clear later, we also take the time derivative of this piecewise solution (17):
@@
One first observation from Eqs. (17) and (18) is that the piecewise solution (17) and also its derivative (18) are proportional to the perturbative force $F$. Thus, in the absence of a perturbative force, the particle stays at rest at the initial position $x(0)=0$ (in the deterministic case).
@@
Second, Eqs. (17) and (18) directly reveal the role of the delay time, $\tau$. Indeed, in the first interval $[0,\tau]$, the system behaves as if there was no delay at all. This changes when we consider the regime $t\geq\tau$: While the first term encodes the systems instantaneous reaction to the applied perturbation, the second and third term can be understood as the time-delayed response. Those terms can lead to a significant change in dynamical behavior, where for each specific $\tau$ the first change in dynamical behavior occurs immediately at $t=\tau$. This means (given a zero history) that from this point on memory effects become important.
@@
Figure 8 presents corresponding exemplary numerical data for Eqs.(17) and (18) where we perturb the system by applying a constant force to the particle starting at $t=0$ for a duration $T_{\mathrm{perturb}}$, which has to be at least longer than one delay time (i.e., $\tau<T_{\mathrm{perturb}}$) and observe the system. The observation continues after the perturbation is set to zero at $t=T_{\mathrm{perturb}}$ for a further duration of $T_{\mathrm{perturb}}$ (Fig. 8c). We observe that after applying the perturbation but also after the perturbation is set to zero, there is a characteristic change in dynamical behavior after one $\tau$ [indicated by color coded vertical dotted lines in Fig. 8a and b]. This change consists in ‘kinks’ and other deviations from a system without delay. We also find that the change is most distinct in $\dot{x}$ (Fig. 8b), even though this is not immediately obvious from the analytical expression (18). Finally, we emphasize that such characteristic response is not limited to linear systems but carries over to more complex nonlinear instantaneous and delayed forces; see Fig. A3 (appendix) for exemplary data.
@@
But how can this knowledge be exploited in a noisy Brownian system with time delay? Does the characteristic response carry over to noisy systems? To this end, in Fig. 8d we show how a linear stochastic system based on feedback forces as in Eq. (2) reacts to a perturbation by a constant external force (Fig. 8e) following a step protocol, where we consider the discretized derivative of positional data. We observe that also in the noisy data the characteristic response of a system with time delay (black lines represent the deterministic response) becomes apparent (Fig. 8d).
@@
In summary, by looking at Fig. 8 and Fig. A3, we notice that the characteristic response of all observed systems herein to an external perturbation is unique to the specific functional form of the forces involved. In particular, it is also present and visible in systems with nonlinear delayed and instantaneous forces. We emphasize that the sudden onset of a sufficiently strong perturbing force is important in order to cause a sufficiently strong reaction compared to noise in stochastic systems. Crucially, and in contrast to the PSD-based approach to inference, instead of observing the trajectory for a long time, the necessary observation duration only has to be of the order of the time delay itself.
@@
Now the question remains, how this feature of systems with time delay can be used to infer the delay time, specifically in noisy Brownian systems. Depending on the parameters, it can be challenging to visually distinguish the characteristic response from noise, despite an average over noise realizations. To this end, we propose the use of a neural network to classify trajectories, or rather stochastic time series, in terms of the underlying delay time. This technique we introduce in the following subsection.
@@ V.2 Convolutional neural network–time series classification
Here, we use this idea to classify one-dimensional time series data (where the time series was produced by application of the perturbation method described above in Subsec. V.1) in terms of the underlying delay time. We make use of the characteristic response of a system with time delay, that is, we assume that the $\tau$-dependent change of dynamical behavior within the time series is a feature the neural network can be trained on. We focus on the discretized derivative of the positional data as shown in Fig. 8d. In addition to the more distinct response compared to that seen in $x(t)$ (see Fig 8a and b), this quantity has the advantage that standardization and normalization (see Appendix A.4) is significantly better achievable compared to pure trajectory data.
@@
Figure 9 shows a simplified network architecture for the classification task. It consists of an input layer which receives the data obtained from the perturbation method in terms of the discretized derivative. Several hidden convolutional and pooling layers follow. The weights and biases associated to the connections within this part of the network are modified in training to minimize an appropriate loss function. Finally, the hidden layers are connected to the output layer, where a softmax activation function is used such that the values of the output layer represent a discrete probability distribution $P(\tau)$ (see rightmost part of Fig 9) over an array of $\tau_{\mathrm{node}}$, i.e., the delay times used in training. This means that each node of the output layer represents one specific delay time $\tau_{\mathrm{node}}$ corresponding to one of the delay times the network was trained on, and the node can assume probability value $P(\tau)$ between zero and one with $\sum_{\tau_{\rm node}}P(\tau)=1.$ Details on network architecture, hyperparameter tuning, and training are given in Appendix A.5. For more information on how to train, test and predict parameters using the convolution neural networks, see Refs. [44, 43].
@@
Figure 9: Visualization of time series classification through convolutional neural networks. Time series data (discretized derivative, see Appendix A.4 for the input data preparation) is fed into the neural network via the input layer. Then, the data goes through a sequence of convolutional and pooling layers. In our specific architecture, before the second last layer the network is flattened to be one-dimensional and subsequently fed to the last hidden layer, a densely connected layer. Finally, a densely connected layer is used as a output layer, where a softmax activation function is employed, in order to return a probability distribution for an array of different $\tau_{\rm node}$. The result is a classification of the time series in terms of delay times.
@@
When training the network, we choose a delay time range as well as the range of the remaining parameters [e.g., for the linear system this means a large number of combinations of the parameters $k_{a}$, $k_{b}$ and $\tau$, each parameter set consists of a tuple $(\tau_{\mathrm{node}},k_{a},k_{b})$, where all are discrete values within the parameter range]. The delay times are represented in the output nodes of the neural network. For each $\tau_{\mathrm{node}}$, the training is then carried out for many combinations of the remaining parameters. This makes the inference results robust, which means that the specific combination of parameters is of minor importance for the inference accuracy of the trained network [44].
@@
Since the goal is to be able to infer time delay from noisy time series data, we train the network on several noise realizations for each of the parameter combinations (of the order of 100 realizations depending on the complexity of the involved delayed and instantaneous forces).

### v2

@@ IV.4 General nonlinear time-delayed feedback
However, we can make use of the observation for the linear system that, at large $\omega$, the minima and maxima are $\Delta\omega\approx 2\pi/\tau$ apart. We hypothesize that oscillations with this oscillation period can also be found in the function $\omega^{2}$PSD of nonlinear systems, i.e., that the signature of time delay carries over from linear to more general nonlinear systems. On this basis, we can design a numerical procedure based on a polynomial fit in order to extract the positions of the minima and maxima in frequency space from numerical data and subsequently compute an approximate inferred delay time (for details of the procedure, see Appendix B.3).
@@
Figure 6: Numerically simulated PSD (red) as a function of $\omega$ for non-linear systems (11), (Panel a) (12) (Panel b), and  (13) (Panel c). In each panel: Polynomial fit (blue) using the quartic splines and automatically extracted positions (cyan circles) of maxima and minima using the roots of the derivative of the polynomial (see Appendix B.3). See Table 1 for the corresponding inference results.
@@
Figure 6 presents exemplary fits of the resulting function $\omega^{2}$PSD including the obtained positions of the minima and maxima, that were used to infer the delay time. The extracted locations of the minima and maxima are in good agreement with the data as shown in Fig. 6, where the extracted locations are marked as circles.
@@
$0.1$ $00\pm 0.002 0.10.1$ $01\pm 0.001 0.10.1$ $0.100\pm 0.005$
$0.15$ $1\pm 0.001 0.150.15$ $2\pm 0.003 0.150.15$ $0.149\pm 0.003$
$0.3$ $03\pm 0.014 0.30.3$ $01\pm 0.009 0.30.3$ $0.300\pm 0.007$
Table 1: Delay times $\tau_{\mathrm{inferred}}$ in units of $\tau_{\mathrm{B}}$ inferred from PSD with one standard deviation as uncertainty: (a) Data corresponding to Eq. (11) with $a=10k_{\mathrm{B}}T/\sigma^{4}$ and $b=c=10k_{\mathrm{B}}T/\sigma^{2}$; (b) Data corresponding to Eq. (12) with $a=10k_{\mathrm{B}}T/\sigma^{4}$, $b=10k_{\mathrm{B}}T/\sigma^{2}$ and $c=20k_{\mathrm{B}}T/\sigma^{3}$; Data corresponding to Eq. (13) with $a=10k_{\mathrm{B}}T/\sigma$, $b=5k_{\mathrm{B}}T/\sigma$ and $c=20k_{\mathrm{B}}T/\sigma$.
Overall, we find robust and accurate inference results for various delay times with an uncertainty of the order of one percent or less. The accuracy of the inference is similar for the different linear and nonlinear feedback and instantaneous forces. Crucially, for the described procedure, knowing the functional form of the forces (instantaneous and delayed) is not required. Therefore, this procedure can be applied to a wide range of systems with discrete time delay. Furthermore, the procedure is not limited to Gaussian white noise systems only. In Appendix D we present numerical results for systems with exponentially correlated noise (see Fig. D5 and Table 3).
@@ V Perturbation and deep learning-based approach – Methodology
Up to this point, we have demonstrated that the power spectral density (PSD) features a characteristic signature of time delay for a broad class of feedback forces (linear and nonlinear) in stochastic systems. This allowed us to infer time delay even if the exact functional form of the underlying forces is unknown. As a drawback, this methodology requires a large number of long trajectories. This can be a challenge in experiments, where one may only have fewer and shorter trajectories at ones disposal.
@@
To circumvent this issue, we here propose an alternative procedure to infer time delay. This procedure is based on the observation of the response of the system under the influence of a perturbative force. Broadly speaking, for a delay-driven system subjected to a perturbation, we expect the system’s response to occur not instantaneously but rather with some delay, as we shall see explicitly in the following Subsec. V.1. Assuming the perturbation to be a force that is applied instantaneously and that is sufficiently strong (i.e., strong compared to noise, feedback forces and instantaneous forces), we expect a distinct delayed reaction, a characteristic response, as we explain in detail in the following.
@@ V.1 Characteristic response of a system with time delay to a perturbation
To develop an analytical understanding of the characteristics of the response, we take one step back and consider a deterministic system (i.e., the limit $T\to 0$) with linear time-delayed feedback force and an additional perturbative force $F_{\mathrm{perturb}}(t)$. Its equation of motion reads
@@
Intuitively, in the limit $k_{a}=k_{b}>0$ (attractive harmonic), the linear feedback force can be understood as the consequence of a delayed harmonic potential centered at $x(t-\tau)$. Figure 7 shows its schematic representation. We assume that the particle was at rest at $x(t)=0$ in the time interval $t\in[-\tau,0]$. We then apply a perturbative force, $F_{\rm perturb}$, starting at $t=0$ for a duration $T_{\mathrm{perturb}}>\tau$. The particle will first be confined to a harmonic potential centered at $x(t-\tau)=0$. Following dynamics (14) after $t=\tau$, the delayed position $x(t-\tau)$ (and hence the location of the harmonic trap) will also start to change, resulting in a change in dynamical behavior.
@@
Figure 8: Effect of a perturbation on deterministic (a-c) and stochastic systems (d-e). Panels (a-c): The particle subject to attractive harmonic time-delayed feedback and external perturbation (14). Red dashed lines: Analytical solutions for position (16) [panel (a)] and velocity (17) [panel (b)] up to $t=2\tau$ for $\tau=0.1$ and up to the end of the perturbation for $\tau=0.15$. Dash-dotted vertical line indicates end of perturbation; dashed grey vertical lines represent the beginning and end of observation; vertical dotted lines (color-coded) serve as a guide for the eye at one delay time $\tau$ after the onset, and then, after the end of the perturbation. Panel (d) shows the discretized ‘velocity’ $v_{x,\mathrm{norm}}$ after noise averaging over $25$ realizations, applying a simple moving average over three time steps and normalization with the maximum velocity $v_{\mathrm{max}}$ as a function of time during the observation window; the deterministic velocity (black, from panel b) is shown as a reference. Dotted vertical lines serve as a guide for the eye at $t=\tau$ and $t=T_{\mathrm{perturb}}+\tau$. Panels (c and e): The perturbing force $F_{\mathrm{perturb}}$ as a function of time. The system is perturbed for a duration of $T_{\mathrm{perturb}}=0.25>\tau$ (a-c) and $T_{\mathrm{perturb}}=0.25\tau_{B}>\tau$ (d-e), and the observation continues for another $T_{\mathrm{perturb}}$.
@@
To make analytical progress, we now assume a perturbation protocol given by a step function, i.e., $F_{\mathrm{perturb}}(t)=F\Theta(t)$, where $\Theta(t)$ is the Heaviside function. We calculate the dynamics resulting from Eq. (14) using the so-called method of steps [78]. This method can be used to obtain an analytical solution in intervals of one delay time iteratively, provided that the history function $\Phi(t)$ containing the position of the particles in the previous time interval is known, i.e., $\Phi(t)=x(t)$ for $t\in[t_{0}-\tau,t_{0}]$. Formally, the piecewise solution as a function of time during the interval $t_{0},t_{0}+\tau]$ reads
$x_{[t_{0},t_{0}+\tau]}(t)=$ $\Phi(t_{0})e^{-\frac{k_{a}}{\gamma}(t-t_{0})}+\frac{k_{b}}{\gamma}\int_{t_{0}}^{t}~dt^{\prime}\Phi(t^{\prime}-\tau)e^{-\frac{k_{a}}{\gamma}(t-t^{\prime})}+\frac{F}{\gamma}\int_{t_{0}}^{t}~dt^{\prime}~\Theta(t^{\prime})e^{-\frac{k_{a}}{\gamma}(t-t^{\prime})}\ .$ (15)
@@
Similarly, the particle’s position $x(t)$ can be evaluated for $t\geq t_{0}+\tau$ iteratively. However, for our analytical considerations, we focus on the dynamical behavior during the first two delay times $\tau$ after switching on the perturbation at $t=0$.
@@
We assume that the initial history function represents a particle at rest in position $x=0$, i.e., we choose $\Phi(t)=x(t)=0$ for $t\in[-\tau,0]$ (which can easily be realized by pinning the particle). The analytical expressions for the position as a function of time during the time interval $\tau,2\tau]$ following Eq. (15) then reads
@@
For reasons that will become clear later, we also take the time derivative of this piecewise solution (16), yielding
@@
One first observation from Eqs. (16) and (17) is that the piecewise solution (16) and also its derivative (17) are proportional to the perturbative force $F$. Thus, in the absence of a perturbative force, the particle stays at rest at the initial position $x(0)=0$, as expected in the deterministic case.
@@
Second, Eqs. (16) and (17) directly reveal the role of the delay time, $\tau$. Indeed, in the first interval $0,\tau]$, the system behaves as if there was no delay at all. This changes when we consider the regime $t\geq\tau$: While the first term encodes the systems instantaneous reaction to the applied perturbation, the second and third term can be understood as the time-delayed response. Those terms can lead to a significant change in dynamical behavior, where for each specific $\tau$ the first change in dynamical behavior occurs immediately at $t=\tau$. This means (given a zero history) that from this point on memory effects become important.
@@
Figure 8 presents corresponding exemplary numerical data including the analytical results from Eqs. (16) and (17). We perturb the system by applying a constant force to the particle starting at $t=0$ for a duration $T_{\mathrm{perturb}}$, which has to be, at least, longer than one delay time (i.e., $\tau<T_{\mathrm{perturb}}$) and observe the system. The observation continues after the perturbation is set to zero at $t=T_{\mathrm{perturb}}$ for a further duration of $T_{\mathrm{perturb}}$ (Fig. 8c). We see from Fig. 8 that after applying the perturbation, but also after the perturbation is set to zero, there is a characteristic change in dynamical behavior after one $\tau$ [indicated by color coded vertical dotted lines in Fig. 8a and b]. This change consists in ‘kinks’ and other deviations from a system without delay. We also find that the change is most distinct in $\dot{x}$ (Fig. 8b), even though this is not immediately obvious from the analytical expression (17). Importantly, such a characteristic response is not limited to linear systems but carries over to more complex nonlinear instantaneous and delayed forces; see Fig. B3 (Appendix) for exemplary data.
@@
But how can this knowledge be exploited in a noisy Brownian system with time delay? Does the characteristic response described before carry over to noisy systems? To this end, in Fig. 8d we show how a linear stochastic system based on feedback forces as in Eq. (2) reacts to a perturbation by a constant external force (Fig. 8e) following a step protocol, where we consider the discretized derivative of positional data. We observe that also in the noisy data the characteristic response of a system with time delay (black lines represent the deterministic response) becomes apparent (Fig. 8d).
@@
In summary, inspecting Fig. 8 and Fig. B3, we notice that all observed systems herein show a characteristic response whose details are unique to the specific functional form of the forces involved. In particular, a characteristic response is also present and visible in systems with nonlinear delayed and instantaneous forces. We emphasize that it is important to have a sudden onset of a sufficiently strong perturbing force to cause a sufficiently strong reaction, compared to noise in stochastic systems. Crucially, and in contrast to the PSD-based approach to inference, instead of observing the trajectory for a long time, the necessary observation duration only has to be of the order of the time delay itself.
@@
Now the question remains, how this feature of systems with time delay can be used to infer the delay time, specifically in noisy Brownian systems. Depending on the parameters, it can be challenging to visually distinguish the characteristic response from noise, despite an average over noise realizations. (In Appendix B.7, Fig. B4, trajectory data for nonlinear systems are presented to give the reader some intuition concerning the complexity of responses and the obfuscation of the change in dynamical behavior through thermal noise.) To this end, we propose the use of a neural network to classify trajectories, or rather stochastic time series, in terms of the underlying delay time. We introduce this technique in the following subsection.
@@ V.2 Convolutional neural network – time series classification
Here, we use this idea to classify one-dimensional time series data (where the time series was produced by application of the perturbation method described above in Subsec. V.1 and in Appendix B.4) in terms of the underlying delay time. We make use of the characteristic response of a system with time delay, that is, we assume that the $\tau$-dependent change of dynamical behavior within the time series is a feature the neural network can be trained on. We focus on the discretized derivative of the positional data as shown in Fig. 8d. In addition to the more distinct response compared to that seen in $x(t)$ (see Fig 8a and b), this quantity has the advantage that standardization and normalization (see Appendix B.4) is significantly better achievable compared to pure trajectory data.
@@
Figure 9 depicts schematically a simplified network architecture for the classification task. It consists of an input layer which receives the data obtained from the perturbation method in terms of the discretized derivative. Several hidden convolutional and pooling layers follow. The weights and biases associated to the connections within this part of the network are modified in training to minimize an appropriate loss function. Finally, the hidden layers are connected to the output layer, where a softmax activation function is used such that the values of the output layer represent a discrete probability distribution $P(\tau)$ (see rightmost part of Fig 9) over an array of $\tau_{\mathrm{node}}$, i.e., the delay times used in training. In other words, each node of the output layer represents one specific delay time $\tau_{\mathrm{node}}$ corresponding to one of the delay times the network was trained on. The node can assume probability values $P(\tau)$ between zero and one with $\sum_{\tau_{\rm node}}P(\tau)=1.$ Details on network architecture, hyperparameter tuning, and training are given in Appendix B.5. For more information on how to train, test and predict parameters using the convolution neural networks, see Refs. [56, 55].
@@
Figure 9: Visualization of time series classification through convolutional neural networks. Time series data (discretized derivative, see Appendix B.4 for the input data preparation) is fed into the neural network via the input layer. Then, the data goes through a sequence of convolutional and pooling layers. In our specific architecture, before the second last layer the network is flattened to be one-dimensional and subsequently fed to the last hidden layer, a densely connected layer. Finally, a densely connected layer is used as a output layer, where a softmax activation function is employed, in order to return a probability distribution for an array of different $\tau_{\rm node}$. The result is a classification of the time series in terms of delay times.
@@
When training the network, we choose a range of delay times as well as a range of remaining parameters. For the linear system this implies a large number of combinations of the parameters $k_{a}$, $k_{b}$ and $\tau$, i.e., each parameter set consists of a tuple $(\tau_{\mathrm{node}},k_{a},k_{b})$, where all are discrete values within the parameter range. The delay times are represented in the output nodes of the neural network. For each $\tau_{\mathrm{node}}$, the training is then carried out for many combinations of the remaining parameters. This makes the inference results robust, with the consequence that the specific combination of parameters is of minor importance for the inference accuracy of the trained network [56].
@@
Since the goal is to be able to infer time delay from noisy time series data, we train the network on several noise realizations for each of the parameter combinations. Specifically, we employ of the order of 100 realizations depending on the complexity of the involved delayed and instantaneous forces.